### v1

# COMMUNITYNOTES: A Dataset for Exploring the Helpfulness of Fact-Checking Explanations

Abstract
Fact-checking on major platforms, such as X, Meta, and TikTok, is shifting from expert-driven verification to a community-based setup, where users contribute explanatory notes to clarify why a post might be misleading. An important challenge here is determining whether an explanation is helpful for understanding real-world claims and the reasons why, which remains largely underexplored in prior research. In practice, most community notes remain unpublished due to slow community annotation, and the reasons for helpfulness lack clear definitions. To bridge these gaps, we introduce the task of predicting both the helpfulness of explanatory notes and the reason for this. We present COMMUNITYNOTES, a large-scale multilingual dataset of 104k posts with user-provided notes and helpfulness labels. We further propose a framework that automatically generates and improves reason definitions via automatic prompt optimization, and integrate them into prediction. Our experiments show that the optimized definitions can improve both helpfulness and reason prediction. Finally, we show that the helpfulness information are beneficial for existing fact-checking systems. The code and data are available at: https://github.com/ruixing76/Helpfulness-FCExp.
COMMUNITYNOTES: A Dataset for Exploring the Helpfulness of Fact-Checking Explanations
Rui Xing1,2 Preslav Nakov2 Timothy Baldwin1,2 Jey Han Lau1 1The University of Melbourne, 2MBZUAI ruixing@student.unimelb.edu.au,preslav.nakov@mbzuai.ac.ae tb@ldwin.net, jeyhan.lau@gmail.com
Figure 1: An example of X Community Notes. The user-generated note appears as “Readers added context.”
1 Introduction
Beyond simple verification, fact-checking systems should provide explanations that clarify why a claim is misleading (Warren et al., 2025). Explanations not only strengthen user trust, but also mitigate pitfalls of fact-checking, e.g., backfire, where the false belief is reinforced rather than hindered (Lewandowsky et al., 2012; Guo et al., 2022; Xing et al., 2025). Typically, fact-checking explanations are produced by expert organizations. Yet the ecosystem is currently shifting, and platforms are increasingly experimenting with community-based fact-checking, where contributors collaboratively provide explanatory notes on claims (Augenstein et al., 2025). This crowdsourcing process distributes the labor of verification and offers diverse perspectives from the community. A representative example is Community Notes on X (formerly Twitter), which allows users to attach notes explaining potentially misleading posts (Figure˜1). [^1] https://communitynotes.x.com/guide/
Critically, for a note to become visible or formally published, it must receive sufficient amount of ratings from a diverse set of users (De et al., 2025). A key challenge in that respect is to decide whether an explanation is helpful for understanding real-world claims and why, which is underexplored in previous research on fact-checking explanations (Atanasova et al., 2020; Guo et al., 2022; Russo et al., 2023; Eldifrawi et al., 2024). This process is also slow on Community Notes: notes can take from hours to days to appear, and the vast majority (over 90%) never become visible, despite the invested contributors efforts. Moreover, while the platform provides reasons for judging notes as helpful or unhelpful, these reasons lack clear definitions, leaving both raters and contributors uncertain about the evaluation criteria (De et al., 2025; Augenstein et al., 2025). This opacity hinders the efficiency, scalability, and explainability of crowd-sourced fact-checking.
Motivated by the above, here we explore the idea of simplifying and making more explicit the helpfulness annotation process. Our contributions are as follows:
  • We introduce the task of predicting the helpfulness and its reasons given a note and the original X post.
  • We present COMMUNITYNOTES, a large-scale explanation helpfulness dataset containing over 104K potentially misleading posts and corresponding user-provided notes.
  • We propose a framework that automatically generates and optimizes reason label definitions via prompt optimization; integrating them to helpfulness/reason predictors to improve their performance.
  • We apply helpfulness prediction to the evidence sufficiency task (Atanasova et al., 2022) and further incorporate it as auxiliary signals into automated fact-checking on CLIMATE-FEVER (Diggelmann et al., 2020), showing its potential to enhance existing fact-checking systems.
2 Related Work
Community-Based Fact-Checking
The proliferation of misinformation has become a growing societal concern, posing significant risks to areas such as public health, science, and politics (Borenstein et al., 2025). To address this challenge, fact-checking has emerged as a central strategy in mitigating the spread of false or misleading information (Guo et al., 2022).
Traditionally, fact-checking has been conducted by professional third-party organizations that evaluate the veracity of claims through expert analysis. However, recent years have witnessed a paradigm shift toward more decentralized, community-based approaches (Renault et al., 2024; Augenstein et al., 2025). Major social media platforms have increasingly embraced this model. X’s Community Notes has become the most prominent example, enabling users to collaboratively annotate and assess the accuracy of online content. Following this trend, Meta recently terminated its partnerships with professional fact-checkers in favor of a similar community-driven system (Catalanello and Sanders, 2025), while TikTok has also introduced a comparable feature (TikTok, 2025).
Empirical evidence suggests that community-based fact-checking can be highly effective in countering misinformation. Studies have shown that the presence of a Community Note can reduce the rate of retweets by nearly half (Renault et al., 2024), while displaying such notes on misleading posts has been found to decrease their spread by an average of 61% (Chuai et al., 2024). Moreover, authors of posts flagged by community notes are up to 80% more likely to delete their content, and such posts generally exhibit markedly lower virality (Drolsbach and Pröllochs, 2023; Renault et al., 2024). However, Community Notes often requires a long time to gather sufficient contributor ratings to determine helpfulness, and the absence of reason definitions further undermines the trustworthiness of the mechanism.
Label Definitions
Recent work has shown that providing models with explicit definitions of labels can lead to substantial improvements in classification. Khatuya et al. (2025) used predefined label descriptions and trained the model to generate these descriptions based on the input text, achieving sizable improvements in multi-label classification. Similarly, Gao et al. (2023) demonstrated that curating a small dictionary of terms for label description can improve zero-shot accuracy by 17-19% absolute, while also increasing the robustness to prompts. Peskine et al. (2023) showed improvements when using label definitions with GPT-3 zero-shot classification on the challenging task of fine-grained conspiracy theory detection.
Automatic Prompt Optimization
Prompt engineering has become essential to many Natural Language Processing (NLP) tasks; yet, manual trial-and-error approaches remain a barrier to effective use. Automatic Prompt Optimization (APO) methods address this challenge by systematically searching for high-performing prompts (Ramnath et al., 2025). These approaches explore the large instruction space using diverse strategies: PromptBreeder evolves task-specific prompts across generations using optimized mutation strategies (Fernando et al., 2024); PromptAgent applies Monte Carlo Tree Search (MCTS) with feedback reflection to guide exploration (Wang et al., 2024); and SCULPT represents prompts as trees, enabling targeted refinements while preserving overall structure—especially useful for long or complex prompts (Kumar et al., 2025).
3 Predicting Whether an Explanation is Helpful and Why
3.1 Task Definition
Given a post $P$ that contains potentially misleading information and a note $N$ that provides an explanation about that post, the task is to predict two types of labels: a helpfulness label $L_{helpful}$, indicating whether the note is helpful in explaining the misleading nature of the post, and a reason label $L_{reason}$, specifying why the note is (un)helpful. The helpfulness prediction is formulated as a binary classification task with the label space $\{Helpful,Unhelpful\}$. The reason prediction is a multi-label classification task, where the predefined label set includes eight reasons for helpful notes and ten reasons for unhelpful ones (See Table˜14 for detailed reason labels).
3.2 Data Collection and Pre-processing
We collected all our data from official Community Notes website between January 2021 and December 2024.222https://communitynotes.x.com/guide/en/under-the-hood/download-data We construct our dataset by combining all publicly available data releases, which include the Notes, Ratings, Note Status History, and User Enrollment tables. Since the ratings data are provided in multiple shards due to their large volume, we first merge all shards and join them with the other components to obtain complete note–rating pairs. We then apply the official note-ranking algorithm to compute the aggregated helpfulness and reason labels for each note. [^3] https://communitynotes.x.com/guide/en/under-the-hood/ranking-notes For helpfulness prediction, we remove entries labeled as NEED_MORE_RATINGS, leaving the helpfulness label space binary—CURRENTLY_RATED_HELPFUL or CURRENTLY_RATED_NOT_HELPFUL. The reason annotation is multi-label, covering 18 predefined categories. Finally, we retrieve tweet content and metadata via the X API and web crawling (Borenstein et al., 2025), linking each post with its corresponding notes to form the final dataset used in our experiments.
Language Train Dev Test Total
English 40,994 5,858 11,717 58,569
Other languages 32,478 4,638 9,281 46,397
Total 73,472 10,496 20,998 104,966
Table 1: Data split statistics for English and other language subsets (updated).
Table˜1 provides the data distribution. Our final dataset contains 104,966 posts, including 58,569 in English and 46,397 in other languages. We perform stratified partitioning with a 7:1:2 ratio for training, development, and testing sets, maintaining separate splits for English and other language data. Henceforth we will refer to this dataset as COMMUNITYNOTES.
3.3 Dataset Statistics
Text Type Mean Median Min Max
Post 57.44 42 0 5,072
Note 87.05 70 1 1,402
Table 2: Token length statistics for claims and notes. Texts are tokenized by OpenAI’s tiktoken v0.11.0.
We inspected multiple aspects of our entire COMMUNITYNOTES dataset. Table˜2 shows the token length distribution for posts and notes correspondingly. The median token number are 42 for posts and 70 for notes, which suggests most posts and notes are short text snippets.
Notes per Post Count (%)
1 84.74%
2 11.84%
3 2.43%
4 0.65%
5 0.18%
6 0.08%
7 0.03%
8 0.01%
$\geq$9 <<0.02%
Table 3: Distribution of the number of notes per post. The majority of posts have only one associated note.
Post Type Count Percentage
All helpful notes 67,558 64.36%
All unhelpful notes 32,851 31.30%
Mixed helpful / unhelpful notes 4,557 4.34%
Table 4: Distribution of posts by helpfulness composition of their associated notes.
In Table˜3, we observe that 84.74% of posts have only one note, 11.84% have two notes, and the remaining posts contain three or more notes. Table˜4 shows that 64.2% of posts have notes that are all rated as helpful, 31.4% have only unhelpful notes, and the remaining 4.4% include notes with mixed ratings. There are eight predefined reasons for a note being rated as helpful and ten reasons for being rated as unhelpful. The most common helpful reasons are helpfulAddressClaim, helpfulImportantContext, and helpfulClear, whereas nothelpfulArgumentativeOrBiased and nothelpfulMissingKeyPoints are the most frequent unhelpful reasons (See detailed reason distribution in Appendix Figure˜4).
Figure˜2 illustrates the language distribution of COMMUNITYNOTES. English is the dominant language, accounting for 57.3% of all notes, followed by Japanese (10.5%) and Spanish (9.9%). This indicates that our final dataset effectively captures the multilingual nature of crowd-sourced explanations.
Figure 2: The languages in the COMMUNITYNOTES: fr-French, es-Spanish, ja-Japanese, pt-Portuguese, de-German, Other: note languages that appear less than 1,000 times.
Beyond analyzing note statistics, we also examine the popularity of the posts targeted by these notes. We use the number of replies, likes, and retweets as metrics (See detailed distributions in Appendix Figure˜5). These posts receive a median of approximately 3k likes, 300 replies, and 700 retweets. This indicates that the posts attracting community notes are highly popular. Given that they are flagged as potentially misleading, this underscores the urgent need to provide helpful explanations for their claims.
3.4 Benchmarking
Experimental Setup
To establish a baseline for understanding the model performance in predicting the helpfulness of the notes, we conducted a comprehensive set of experiments using both Small Language Models (SLMs) (BERT-style encoder-only architectures in our paper) and decoder-only Large Language Models (LLMs). For SLMs, we formulate the task as multi-task classification, where the model is jointly trained to predict (1) a binary helpfulness label and (2) a multi-label reason classification across 18 predefined categories. The model input is constructed as a concatenated sequence in the format “[CLS] Claim: {claim} [SEP] Note: {note_text}” and subsequently passed through two separate classification heads. For LLMs, we fine-tune models in the same multi-task way (See detailed prompt in Appendix Table˜14). We partitioned our experiments by model and language setting. For English, we experimented with SLMs including BERT, ModernBERT, RoBERTa and DeBERTa (both base and large). We also evaluated the performance of LLMs including Llama3.1-8B and Mistral-7B, under both zero-shot and LoRA fine-tuning conditions (Hu et al., 2022). For the multilingual part, we used models specifically trained on multilingual data, including BERT-base-multilingual (mBERT) and XLM-RoBERTa (base and large).
Language Model Helpfulness F1 Reason P Reasons R Reason F1
English BERT-base 0.866 0.549 0.742 0.631
BERT-large 0.874 0.553 0.758 0.640
RoBERTa-base 0.876 0.546 0.783 0.643
RoBERTa-large 0.890 0.581 0.772 0.663
ModernBERT-base 0.871 0.545 0.755 0.633
ModernBERT-large 0.888 0.576 0.752 0.652
DeBERTa-base 0.886 0.541 0.810 0.650
DeBERTa-large 0.896 0.589 0.759 0.665
Llama3.1-8B-ins 0.919 0.640 0.640 0.640
Mistral-7B-ins 0.920 0.620 0.620 0.620
Other mBERT 0.873 0.562 0.799 0.659
XLM-RoBERTa-base 0.882 0.584 0.798 0.674
XLM-RoBERTa-large 0.894 0.603 0.815 0.693
Llama3.1-8B-ins 0.928 0.653 0.652 0.653
Mistral-7B-ins 0.926 0.647 0.647 0.647
Table 5: Benchmark results for helpfulness and reasons prediction. Helpfulness means the binary classification of notes helpfulness. Reason means multi-label classification of helpfulness reasons.
Results and Analysis
The models generally perform well on the helpfulness prediction task, with most of them achieving F1 scores beyond 0.88. This suggests that distinguishing between helpful and non-helpful notes is a relatively easy task. In contrast, the reasons prediction task is much more challenging. This is showed by the much lower F1 scores, particularly the F1 scores which consistently fall below 0.70.
Comparing SLMs to LLMs: The table shows that LLMs like Llama3.1-8B and Mistral-7B outperform smaller SLMs on Helpfulness prediction. The Mistral-7B model achieves the highest F1 Score (0.920), demonstrating the superior capabilities of larger, more advanced models. However, the LLMs perform worse than SLMs in reason prediction, with their F1 scores being comparable to the best-performing SLMs like DeBERTa-large. Within the SLM families (BERT, RoBERTa, DeBERTa), larger models consistently outperform their base-sized counterparts. For example, RoBERTa-large achieves better scores than RoBERTa-base across all metrics. This highlights the importance of model capacity for capturing the nuances required for this complex task. The table also shows that DeBERTa models generally perform at or near the top among the SLMs, particularly for the reasons task with the highest F1 Score of 0.665 among all SLMs.
The multilingual part show a similar trend: larger models achieve strong performance in Helpfulness prediction while smaller models like mBert and XLM-RoBERTa exhibit better performance on the reason prediction. Among all models, XLM-RoBERTa-large achieves the highest Reason F1 of 0.693. This indicates that these models are effective at generalizing across different languages.
Figure 3: Our framework for automatic reason definition generation and optimization.
4 Enhancing Helpfulness Prediction with Reason Definitions
Preliminary analysis shows that predicting why a particular explanation is (un)helpful is often subjective across instances, largely due to the lack of unified or official definitions. This subjectivity poses a barrier to model performance and limits the explainability of community-based fact-checking platforms. Prior work has demonstrated that providing clear label definitions can significantly improves the classification performance (Gao et al., 2023; Peskine et al., 2023; Khatuya et al., 2025). Given the large number of notes with annotated reason available, we aim to automatically generate and optimize reason label definitions based on data. To this end, we introduce a pipeline that leverages Automatic Prompt Optimization (APO) to generate and improve helpfulness reason definitions, which are then integrated with notes for reason prediction. Since the majority of notes and definitions are written in English, we focus on English data for the remainder of our experiments. Figure˜3 illustrates the overall pipeline.
4.1 Automatically Optimizing Reason Definition
We generate and optimize reason definitions following the first two steps in Figure˜3: (1) Seed reason definition generation and (2) Reason definition optimization. In the first step, we follow Peskine et al. (2023) and randomly sample 40 instances for each category. We then prompt the GPT-4o model to generate candidate definitions (see Table˜13 in Appendix for detailed prompts). [^4] Since a single sample may have multiple reason labels, we perform an additional check to ensure that each sample appears only within the same category to encourage diversity.
In the second step, we employ PromptAgent (Wang et al., 2024), a Monte Carlo Tree Search (MCTS)-based automatic prompt optimization framework, to refine the initial definitions. PromptAgent adopts an iterative refinement process that emulates how human experts strategically craft superior prompts, implemented through an MCTS-based planning mechanism. Each refined prompt corresponds to a new state in the search space, obtained by the refiner agent model through feedback collected from a feedback agent model. PromptAgent relies on the MCTS state–action value function to estimate the potential reward of each state. MCTS iteratively performs four key operations—selection, expansion, simulation, and backpropagation—until a predefined number of iterations is reached. The highest-reward trajectory is then selected to produce the final refined prompt. We initialize the process with the seed definitions generated in step (1) and run PromptAgent using a definition-refinement feedback loop.
Model Helpfulness F1 Reason F1
Llama3.1-8B-ins 0.636 0.126
Llama3.1-8B-ins-seed 0.783 0.304
Llama3.1-8B-ins-opt 0.785 0.344
Mistral-7B-ins 0.761 0.135
Mistral-7B-ins-seed 0.776 0.257
Mistral-7B-ins-opt 0.786 0.324
Table 6: Zero-shot prediction performance for LLMs w/o adding reason definitions on our COMMUNITYNOTES test set. No suffix means the vanilla model prediction performance. seed means definitions are generated by GPT-4o on randomly sampled cases. opt means that the definitions are further optimized by PromptAgent.
To evaluate the impact of incorporating generated definitions, we perform zero-shot prompting with Llama3.1-8B-Instruct and Mistral-7B-Instruct for both helpfulness and reason prediction. Table˜6 reports results for the vanilla setting, as well as when incorporating seed and optimized definitions. We observe substantial F1 improvements in both helpfulness and reason prediction when adding seed definitions (on average around 15%), with further gains after definition optimization. For Llama3.1-8B-Instruct, both seed and optimized definitions yield large improvements. Although Mistral-7B-Instruct starts stronger than Llama3.1-8B-Instruct in the vanilla setting, it still benefits notably from incorporating definitions, particularly in reason-level F1. These results suggest that incorporating reason definition information substantially enhances model performance—not only in predicting whether a given explanation note is helpful, but also in identifying the underlying reason with respect to the claim (See seed definition in Table˜17 and optimized reason definition in Appendix LABEL:tab:optimized_def).
4.2 Training with Reason Definition Fusion
After obtaining the optimized helpfulness reason definitions, we proceed to the third step in Figure˜3, where we explore how to incorporate this information into the predictor’s training process. The core idea is that, for a given note associated with a claim, the predictor should attend more strongly to the relevant helpfulness reasons. Since SLM performs better in reason prediction, we use SLM as our predictors in the experiments.
We first obtain note embeddings $E_{note}$ from the original predictors and reason embeddings $E_{reason}$ from the optimized reason definitions using LLM2Vec (BehnamGhader et al., 2024), an unsupervised LLM-based embedding model. [^5] Specifically, we use LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised. We then apply a Multi-Head Attention (MHA) mechanism, using $E_{note}$ as the query (Q) and $E_{reason}$ as the key (K) and value (V), to fuse the reason information into the note representation. Finally, we adopt a multi-task learning objective that jointly optimizes helpfulness prediction and multi-label reason classification.
Table˜7 shows the impact of integrating the optimized reason definitions into the base models using the MHA module. Across all predictors, incorporating reason definitions via MHA consistently improves both helpfulness and reason F1 scores. For instance, RoBERTa-large-MHA achieves a Helpfulness F1 of 0.899 and a Reason F1 of 0.669, outperforming its vanilla counterpart (0.890 and 0.663, respectively). Similar trends hold for ModernBERT and DeBERTa, with DeBERTa-large-MHA achieving the highest Reason F1 score of 0.677. These results demonstrate that attention-based fusion of reason definitions effectively enhances SLMs, improving their ability to predict both whether a note is helpful and the underlying reason.
Model Helpfulness F1 Reason F1
RoBERTa-base 0.876 0.643
RoBERTa-base-MHA 0.887 0.652
RoBERTa-large 0.890 0.663
RoBERTa-large-MHA 0.899 0.672
ModernBERT-base 0.871 0.633
ModernBERT-base-MHA 0.874 0.644
ModernBERT-large 0.888 0.652
ModernBERT-large-MHA 0.898 0.662
DeBERTa-base 0.886 0.650
DeBERTa-base-MHA 0.887 0.668
DeBERTa-large 0.896 0.665
DeBERTa-large-MHA 0.896 0.677
Table 7: Performance comparison of base models versus fusing optimized reason definition with the Multi-Head Attention module.
5 Generalization of the Explanation Helpfulness
In this section, we assess the generalization and the utility of our helpfulness predictors beyond community notes. A key challenge is that most existing datasets lack explicit annotations for explanation of helpfulness. Therefore, we explore two key scenarios in automated fact-checking. First, we evaluate whether helpfulness prediction can generalize to evidence sufficiency prediction, which is an important subtask in fact-checking. Second, we explore the effect of incorporating evidence helpfulness in automated fact-checking.
5.1 Generalization to Evidence Sufficiency
Evidence sufficiency is a task that decides whether the information in the evidence is sufficient for predicting the veracity of a claim. Automated fact-checking models can make veracity predictions only when there is enough evidence (Atanasova et al., 2022). In this scenario, we would like to explore whether our helpfulness predictors can predict the sufficiency of an evidence (since a helpful evidence must be “sufficient”). We adopted Sufficient Fact (Atanasova et al., 2022), a diagnostic fact-checking dataset to detect when evidence with omitted information is (in)sufficient. To adapt our prediction, we treat sufficient evidence (EI-ENOUGH INFO) as helpful and insufficient evidence (NEI-NOT ENOUGH INFO) as unhelpful.
Predictors NEI P NEI R NEI F1
RoBERTa-base-FT 0.869 0.920 0.894
RoBERTa-base-notes 0.742 0.521 0.612
RoBERTa-large-FT 0.882 0.935 0.908
RoBERTa-large-notes 0.731 0.676 0.702
ModernBERT-base-FT 0.845 0.843 0.844
ModernBERT-base-notes 0.770 0.779 0.774
ModernBERT-large-FT 0.902 0.906 0.904
ModernBERT-large-notes 0.808 0.807 0.808
DeBERTa-base-FT 0.912 0.855 0.882
DeBERTa-base-notes 0.746 0.553 0.635
DeBERTa-large-FT 0.919 0.890 0.905
DeBERTa-large-notes 0.736 0.774 0.755
Table 8: Predictor generalization performance on evidence sufficiency task. notes stands for models for original communitynotes predictors. FT means FineTuned models on evidence sufficiency task. NEI means Not Enough Information label.
Table˜8 shows the model performance on the evidence sufficiency task, focusing on the NEI label. As expected, fine-tuned (FT) models outperform the original notes predictors, with RoBERTa-large-FT achieving the best NEI F1 of 0.908. Nevertheless, the notes models show competitive performance: for example, ModernBERT-large-notes achieves a NEI F1 of 0.808, only about 10% lower than its FT counterpart. Across all predictors, the top predicted reason is notHelpfulMissingKeyPoints, which closely corresponds to the NEI label, as both signify the absence of critical information.
This indicates that helpfulness prediction carries transferable signal for evidence sufficiency. Still, helpfulness and sufficiency are not the same: even sufficient evidence could be predicted as unhelpful for other reasons such as notHelpfulArgumentativeOrBiased, which may account for the observed performance gap.
5.2 Incorporating Evidence Helpfulness in Automated Fact Checking
In this part, we further investigate whether predicted evidence helpfulness can enhance automated fact-checking. Helpfulness scores provide auxiliary signals for assessing evidence quality, with the potential to help models prioritize informative evidence during prediction. We adopted CLIMATE-FEVER (Diggelmann et al., 2020), a real-world dataset for verifying climate change–related claims. Each claim is associated with multiple evidence sentences, to which we apply our predictors to assign helpfulness scores and reasons. To perform fact-checking, we use Llama3.1-8B-Instruct to perform fact-checking (See detailed prompt in Appendix LABEL:tab:fc_prompt) prompting the model with both the claim and its associated evidence.
We compare the F1 with and without the helpfulness information of evidence and report results in Table˜9. We can see that, in general, incorporating helpfulness improves fact-checking performance with ModernBERT-base achieved an accuracy of 0.537. ModernBERT and RoBERTa-base improve fact-checking performance by around 2%. These findings indicate that providing auxiliary evidence helpfulness and reasons can benefit automated fact-checking. While ModernBERT demonstrates major gains, DeBERTa and RoBERTa yield smaller, non-significant improvements. Interestingly, this pattern mirrors their performance drop in cross-domain evaluation of evidence sufficiency in Section˜5.1 (despite having the best in-domain results in Table˜5). One possible explanation could be that these models have overfitted for the Community Notes domain and as such are less effective in other domains.
Predictors F1 +Helpfulness F1
RoBERTa-base 0.517 0.521
RoBERTa-large 0.519 0.521
ModernBERT-base 0.521 0.535
ModernBERT-large 0.517 0.535
DeBERTa-base 0.517 0.519
DeBERTa-large 0.518 0.519
Table 9: Automated fact-checking performance of Llama3.1-8B-instruct on the CLIMATE-FEVER dataset. F1 stands for Accuracy. +Helpfulness F1 means accuracy incorporating helpfulness information. Numbers in bold represents the gap between +Helpfulness and basic accuracy is statistically significant (p-value <0.05).
6 Conclusion
In this paper, we introduced the novel task of predicting the helpfulness of explanations and their underlying reasons for a given claim. To support this task, we constructed COMMUNITYNOTES, a large-scale post–note helpfulness dataset based on X’s Community Notes, and established benchmarks with pretrained language models. While existing models perform reasonably well on overall helpfulness, they struggle to identify underlying reasons. To address this, we proposed a framework that automatically generates and optimizes reason definitions, which, when reintegrated, greatly improves reason prediction. We further demonstrated that helpfulness prediction can help predict evidence sufficiency and enhance fact-checking systems.
We hope that our dataset and framework can serve as a foundation for building stronger helpfulness predictors capable of assigning prior labels to real-time community notes, thereby streamlining the community-based fact-checking process. Additionally, the automatically generated reason definitions may enhance the explainability and trustworthiness of fact-checking models and could be extended to improve other classification tasks beyond this domain.
Limitations
While our study provides new insights into the helpfulness of crowd-sourced fact-checking explanations, several limitations remain.
First, the inherently subjective nature of community-based annotations introduces potential noise and inconsistency in the helpfulness and reason labels. Despite our efforts to aggregate ratings through official mechanisms, differences in annotator interpretation and cultural context may still influence the results.
Second, our automated definition generation and optimization pipeline has thus far been evaluated primarily on English. Although our dataset includes multilingual content, we did not systematically validate the quality or transferability of the generated definitions across languages. Future work should explore multilingual adaptation and cross-lingual alignment of reason definitions.
Third, while our experiments show improved model performance when integrating optimized reason definitions, the evaluation is limited to classification metrics. Further human-centered evaluation is necessary to assess whether these improvements translate to greater interpretability and practical usefulness in real-world fact-checking workflows.
Finally, our study focuses on publicly available Community Notes data, which may reflect platform-specific norms and biases. For a broader generalization to other fact-checking ecosystems (e.g., Meta’s or TikTok’s community moderation systems), we need further investigation.
Ethics and Broader Impact
Our work aims to improve the transparency, explainability, and scalability of community-based fact-checking systems. By modeling and predicting the helpfulness of user-generated explanations, we hope to assist both contributors and platforms in identifying notes that are more informative, balanced, and useful for the public discourse. However, several ethical considerations deserve discussion.
Bias and Fairness.
Community-based ratings may reflect social, cultural, or political biases that influence which notes are deemed “helpful.” Our models inevitably inherit such biases from the underlying data. We emphasize that model predictions should not be used to automatically moderate or suppress user content. Instead, they should be interpreted as supporting signals to guide human review.
Responsible Use.
While our framework can aid in prioritizing or surfacing potentially helpful notes, it is not intended to replace human judgment. Automated systems that assign helpfulness or credibility labels at scale could have downstream impacts on content visibility, user reputation, and the dynamics of online discourse. We encourage future work to incorporate fairness-aware and interpretability-centered mechanisms before deployment.
Positive Societal Impact.
Despite these caveats, we believe our dataset and models can contribute positively to combating misinformation by improving the quality, trustworthiness, and accessibility of community fact-checking. By making the definitions of helpfulness explicit and machine-interpretable, we hope to support more transparent, accountable, and participatory approaches to information verification.
Data Collection and Licenses
All data used in this work originate from the publicly available X Community Notes dataset, [^6] https://communitynotes.x.com/guide/en/under-the-hood/download-data which is released by the platform under its open data policy. The dataset consists of user-generated notes, their associated posts, and corresponding helpfulness and reason labels. No personally identifiable information (PII), user handles, or private metadata are included in our version of the dataset. We processed and released only anonymized text and aggregated statistics in compliance with the platform’s terms of service. The dataset is multilingual, with English, Japanese, and Spanish as the most frequent languages. All examples are drawn from publicly visible content. Any future release of derived data or annotations will follow ACL’s data ethics guidelines and include proper documentation describing preprocessing steps, filtering criteria, and licensing terms.
Security Implication
Our study does not introduce direct security risks, but the use of models trained on user-generated content entails potential vulnerabilities. Automated systems that evaluate or rank explanatory notes could be misused for large-scale manipulation—for example, to promote particular narratives or suppress dissenting perspectives. To mitigate such risks, we release only aggregated and anonymized data, and we emphasize that our models are intended for research and analysis, not for automated moderation or deployment without human oversight. Future work should include safeguards against adversarial attacks, bias exploitation, and model inversion risks, particularly when integrating such predictors into live social platforms.
Appendix A Appendix
A.1 Extended Background of X Community Notes
Origin
Community Notes originated from the Birdwatch (Wojcik et al., 2022) program launched by X (formerly Twitter) in 2021.777Introducing Birdwatch, a community-based approach to misinformation, By Keith Coleman, Monday, 25 January 2021 Participants, known as contributors, write “notes” that clarify or contextualize claims made in posts, and other users rate these notes for helpfulness and accuracy. Notes receiving broad agreement across diverse raters are surfaced publicly beneath the corresponding posts. This community-driven system aims to promote transparency and provide multiple perspectives in addressing misinformation on social media (Augenstein et al., 2025). Initially limited to U.S. users, the program gained wider attention in 2022 during major misinformation events such as the Russian invasion of Ukraine and the COVID-19 pandemic. Birdwatch was re-branded as Community Notes and expanded globally in November 2022. As of November 2023, the program had over 133,000 contributors, with notes reportedly viewed tens of millions of times per day, reflecting its growing role in mitigating misinformation online.
Advantages and Challenges
Community Notes introduces a promising new paradigm for fact-checking. It democratizes content moderation by enabling the public to collectively determine whether a note should be added to a post, offering a less intrusive alternative to traditional expert-based fact-checking systems that often rely on explicit warning labels. This community-driven approach also accelerates the detection of misinformation, offering better scalability. However, Community Notes also faces several challenges. The community can still favor popularity over truth, similar to the echo-chamber effects often seen on social media. While its scalability and faster operation are promising, studies show that its performance still lags behind expert-based fact-checking in many cases (Martel and Rand, 2024). In addition, helpful notes often appear too late—after a post has already received significant engagement—reducing their effectiveness in correcting misinformation and changing user beliefs (Augenstein et al., 2025; De et al., 2025).
A.2 COMMUNITYNOTES Pre-processing Details
Raw Data
Community notes data are publicly available on the official website. [^8] https://x.com/i/communitynotes/download-data The data is released in several following separate files:
  • Notes: This table contains all notes and their attributes. Core attributes include tweetId, which identifies the tweet the note is associated with, and classification, indicating whether the tweet is considered misleading. It also includes several attributes describing potential types of misinformation (e.g., misleadingFactualError); however, these fields are not directly related to note ratings and are excluded from this study.
  • Ratings: This table includes all user ratings of notes. Our helpfulness label scheme is derived from this table, where attributes prefixed with helpful- and notHelpful- correspond to the helpfulness reasons used in our COMMUNITYNOTES experiments.
  • Note Status History: This table contains metadata about notes, including their status updates and timestamps. We use the currentStatus attribute to determine each note’s final helpfulness label.
  • User Enrollment: This table provides metadata on each user’s enrollment status in the Community Notes program. Since it does not contain information related to note helpfulness, we exclude it from our experiments.
  • Note Requests: This table records all requests submitted by X users for a Community Note. It was added after our data collection period, so it is not used in this study.
We first construct the full dataset by joining all data files on the noteID column. We then remove empty notes and merge the duplicated helpfulness reasons notHelpfulOpinionSpeculation and notHelpfulOpinionSpeculationOrBias into a single category, notHelpfulOpinionSpeculationOrBias. Finally, we exclude notes labeled as notHelpfulOther, as their reasons are unclear and uninformative. Table˜10 presents data size statistics before and after processing.
Language Before Processing After Processing
Train Dev Test Train Dev Test
English 41,102 5,872 11,744 40,994 5,858 11,717
Other languages 32,520 4,646 9,292 32,478 4,638 9,281
Total 73,622 10,518 21,036 73,472 10,496 20,998
Table 10: Comparison of dataset sizes before and after processing for English and multilingual subsets.
Note Ranking Algorithm
Contributors submit and rate community notes. Community Notes use these ratings to determine note statuses (CURRENTLY_RATED_HELPFUL, CURRENTLY_RATED_NOT_HELPFUL, or NEED_MORE_RATINGS). Note status determines which note is displayed on each of the Community Notes Site. Table˜11 illustrates the entire algorithm adopted by Community Notes. All notes start with NEED_MORE_RATINGS until they receive at least five ratings. The algorithm is based on a Matrix Factorization (MF) model fit on the sparse note-rater matrix which generates a continuous "Note Helpfulness Score" for each note. A note receives CURRENTLY_RATED_HELPFUL status if its score is greater than 0.40. Currently, only "potentially misleading" notes with this status are shown publicly. A note is assigned CURRENTLY_RATED_NOT_HELPFUL if its score is smaller than -0.05 - 0.8 $\times$ $|\text{noteFactorScore}|$, or if its upper confidence bound is smaller than -0.04. Notes with scores between these "Helpful" and "Not Helpful" thresholds remain in the NEED_MORE_RATINGS status. We reproduced the entire algorithm on a 64-core Intel(R) Xeon(R) Gold 6448H CPU with 500 GB of RAM to obtain the aggregated helpfulness reasons for each note in our COMMUNITYNOTES. [^9] https://github.com/twitter/communitynotes
Phase # Step Description
Prescoring 1. Pre-filter Data: Include only raters ($\geq 10$ ratings) & notes ($\geq 5$ ratings). Merge similar ratings.
2. Initial Model Fit: Fit matrix factorization (MF) models for each scorer. Assign intermediate note status based on intercept thresholds.
3. Calculate & Filter Helpfulness: Compute Author/Rater Helpfulness Scores from the initial MF results. Filter out low-helpfulness raters.
4. Refine Helpfulness: Fit a tag-consensus MF model on the helpfulness-filtered data and update all helpfulness scores.
Scoring 1. Load & Refresh Data: Load prescoring outputs. Re-run the pre-filtering step on the newest available data.
2. Re-fit Main Models: Re-fit the main MF models (for all scorers) using the helpfulness-filtered ratings.
3. Fit Diligence Model: Fit the note diligence MF model.
4. Compute Confidence Bounds: Add pseudo-ratings and re-fit models to get upper/lower confidence bounds for note intercepts.
5. Reconcile & Finalize Status: Generate a final status for each note from all scorer results. Stabilize status for notes older than two weeks.
6. Assign Explanation Tags: Assign the top two matching explanation tags. If two such tags are not found, revert status to "Needs More Ratings".
Table 11: Overview of the X Community Notes Note Ranking Algorithm Pipeline
Figure 4: A historgram of the Note Reason label frequencies. COMMUNITYNOTES contains 18 reason categories—8 corresponding to helpful notes and 10 to not helpful notes.
A.3 Extended COMMUNITYNOTES Details
We present COMMUNITYNOTES Helpfulness Reason distribution in Figure˜4 and Tweets Reaction statistics in Figure˜5. Three “helpful” categories account for the vast majority of responses: helpfulClear (52,921), helpfulImportantContext (52,384), and helpfulGoodSources (48,134). These are followed by a distinct middle tier of common “not helpful” reasons, led by notHelpfulNoteNotNeeded (23,427), notHelpfulMissingOrUnreliable (19,411), and notHelpfulOpinionSpeculationOrBias (17,446). The remaining reasons form a long tail with substantially lower frequencies. As shown in Figure˜5, The mean reaction counts are substantial (e.g., $\approx$ 1,200 replies, $\approx$ 2,800 retweets, and $\approx$ 20,000 likes), indicating that posts in the dataset receive significant public attention on average. However, the median (solid green line) for all three metrics is near zero. This large gap between the mean and median confirms the existence of a long tail of highly viral tweets that receive tens of thousands of reactions. The high visibility and broad reach of these popular posts underscore the urgent need for scalable community-based fact-checking.
Figure 5: Boxplot of tweet reactions in COMMUNITYNOTES. We use three metrics for reactions, numebr of likes, reply counts and retweets.
A.4 Prompts
In this section, we provided detailed prompts used in our COMMUNITYNOTES experiments. Table˜17 shows the prompt for initial reason generation. Table˜14 presents the prompt for LLM to predict Helpfulness and Reasons. Table˜15 and Table˜16 show prompts with seed definition and optimized definition for LLM predictions. LABEL:tab:optimized_def is the detailed reason definition optimized using PromptAgent. LABEL:tab:fc_prompt shows the prompt we used for LLM to perform fact checking on CLIMATE-FEVER dataset.
Model Training Epoch Batch Size
BERT-base 19 64
BERT-large 8 32
RoBERTa-base 16 64
RoBERTa-large 10 32
ModernBERT-base 6 32
ModernBERT-large 5 16
DeBERTa-base 10 32
DeBERTa-large 8 16
XLM-RoBERTa-base 20 64
XLM-RoBERTa-large 10 32
mBERT 11 64
Table 12: Training details for SLMs.
A.5 Experimental Details
We report the best training epochs and batch sizes for SLMs in Table˜12. To reduce overfitting, we use smaller batch sizes and fewer epochs for base models, and larger ones for their large counterparts. For LLMs, we apply LoRA fine-tuning with a rank of 16, alpha of 32, and applied to all linear modules. We train with a batch size of 4, gradient accumulation of 1 step, for 3 epochs, using a cosine learning rate scheduler with a warmup ratio of 0.1. The training takes between 32 and 46 GPU hours on 2 NVIDIA RTX6000 Ada Generation GPUs (48GB GPU RAM each). All models are directly adopted from the Hugging Face Hub, including SLMs (BERT  [^10] https://huggingface.co/docs/transformers/en/model_doc/bert, RoBERTa  [^11] https://huggingface.co/docs/transformers/en/model_doc/roberta, DeBERTa  [^12] https://huggingface.co/docs/transformers/en/model_doc/deberta, ModernBERT  [^13] https://huggingface.co/docs/transformers/en/model_doc/modernbert, mBERT  [^14] https://huggingface.co/google-bert/bert-base-multilingual-cased, XLM-RoBERTa  [^15] https://huggingface.co/docs/transformers/en/model_doc/xlm-roberta) and LLMs (Llama3.1-8B  [^16] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Mistral-7B  [^17] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3).
Initial reason generation
You will be given a set of samples, each sample contains CLAIM, their corresponding NOTE to explain the CLAIM. All samples provided are \${helpful_label} in explaining the CLAIM and associated with the same REASON. Your task is to conclude the definition of the REASON. Here are samples: \${samples} The REASON for above samples being \${helpful_label} is \${reason_label}. After checking these samples, the definition of this REASON is:
Table 13: Prompt for initial reason generation
Original Prompt
Given a potentially misleading CLAIM and an associated NOTE, your task is to determine whether the NOTE is helpful in clarifying the CLAIM and identify two reasons from the predefined reason set explaining why it is helpful or not helpful.
The predefined reason set is:’helpfulAddressesClaim’, ’helpfulClear’, ’helpfulEmpathetic’, ’helpfulGoodSources’, ’helpfulImportantContext’, ’helpfulInformative’, ’helpfulUnbiasedLanguage’, ’helpfulUniqueContext’, ’notHelpfulArgumentativeOrBiased’, ’notHelpfulHardToUnderstand’,’notHelpfulIncorrect’, ’notHelpfulIrrelevantSources’, ’notHelpfulMissingKeyPoints’, ’notHelpfulNoteNotNeeded’, ’notHelpfulOffTopic’, ’notHelpfulOpinionSpeculation’, ’notHelpfulOpinionSpeculationOrBias’, ’notHelpfulOther’, ’notHelpfulSourcesMissingOrUnreliable’, ’notHelpfulSpamHarassmentOrAbuse’
Output in the following JSON format only, no extra output: {"helpfulness": helpful or non_helpful,"reasons":"reason1;reason2"} CLAIM: \${claim}
NOTE: \${note}
Answer:
Table 14: Original English prompt for note helpfulness and reason prediction.
Prompt with seed defininion
Given a potentially misleading CLAIM and an associated NOTE, your task is to determine whether the NOTE is helpful in clarifying the CLAIM and identify two reasons from the predefined reason set explaining why it is helpful or not helpful.
Here are reasons and their definitions: \${reason definitions} Output in the following JSON format only, no extra output: {"helpfulness": helpful or non_helpful,"reasons":"reason1;reason2"}
CLAIM: \${claim}
NOTE: \${note}
Answer:
Table 15: English prompt with seed definition for note helpfulness and reason prediction.
Optimized Prompt
For each claim-note pair below, select exactly two reasons (from the provided list) that most accurately explain whether and why the note is helpful or not helpful for understanding or resolving the claim.
High-Level Criteria
1. Prioritizing Central, Accurate Clarification
Assign a “helpful” reason only if the note is factually accurate, well-supported, and directly clarifies, corrects, or provides essential context for the main assertion, a key factual sub-claim, or the foundational evidence of the claim. Peripheral or partially related details are insufficient. If the note’s chief function is to correct a specific fact, figure, identity, date, source, or a significant misattribution or misconception central to the claim, this counts as clarification or correction (use “helpfulClear” and/or “helpfulAddressesClaim”). Never assign any “helpful” reason if the note contains substantive factual errors, misrepresentations, speculation, or bias regarding any key claim point or related evidence—regardless of any attempt to clarify. 2. “Helpful” Reason Prioritization helpfulClear and helpfulAddressesClaim take precedence over other “helpful” reasons when a note offers a direct factual correction or explicit clarification for the claim’s core point or supporting evidence. helpfulGoodSources should only be selected in addition if the note’s correction is fundamentally grounded in clearly cited, authoritative, and faithfully summarized sources (not just the presence of a source). helpfulImportantContext is used when the note supplies background or context without which the claim would be misunderstood or misinterpreted, and this context is not a direct factual correction but provides essential interpretive clarity. Use helpfulUnbiasedLanguage or helpfulEmpathetic only in conjunction with a substantive clarification or correction and if the neutrality or tone of the explanation is materially helpful. When both direct correction and crucial context are present, prefer (b) helpfulClear or (a) helpfulAddressesClaim as primary, paired with (e) helpfulImportantContext as secondary only if context is indispensable and not redundant with the correction. helpfulGoodSources never substitutes for correction—pick it only if the faithful use of sources is a main reason for helpfulness.
3. “NotHelpful” Reason Selection Assign “notHelpfulIncorrect” if there is any inaccuracy, factual misstatement, or misleading claim regarding the main point or evidence. Assign “notHelpfulMissingKeyPoints” if the note avoids or omits addressing the core issue or supporting fact, no matter how detailed its peripheral info. Use “notHelpfulNoteNotNeeded” if the note is trivial, redundant, or supplies information already evident and non-essential for claim comprehension. “notHelpfulSourcesMissingOrUnreliable” or “notHelpfulIrrelevantSources” apply if sources are not credible, are misrepresented, or are irrelevant to the core of the claim. If tone, personal opinion, bias, speculation, or argumentativeness blocks any clarification, use one of the corresponding “notHelpful” reasons that best fits the limitation. For unclear, incomplete, or confusing notes, use “notHelpfulHardToUnderstand.” “notHelpfulOther” and “notHelpfulOffTopic” only if none of the above describe the problem or the note is wholly irrelevant. Reason Definitions:\${Reason Definitions}
Decision Process Checklist
1. Is the note factually accurate and not misleading about any major point or evidence? If no, assign “notHelpfulIncorrect” and another as fits. If yes, proceed. 2. Does the note directly clarify, correct, or critically contextualize the main assertion, a key factual sub-claim, or any central supporting evidence? If yes, select the most specific “helpful” reason(s) per above priority order. If its primary value is sources, include “helpfulGoodSources” only if the sourcing itself is decisive. Do not select “helpfulImportantContext” unless the info is both necessary for accurate interpretation and not primarily a direct correction. If note only offers peripheral detail, trivia, or sidesteps the key issue, use “notHelpfulMissingKeyPoints” and/or “notHelpfulNoteNotNeeded.” 3. Is the note clear, neutral, and respectful in tone? If so in addition to being factually helpful, pair with “helpfulUnbiasedLanguage” or “helpfulEmpathetic” as needed. If tone, speculation, or bias prevents meaningful clarification, pick the corresponding “notHelpful” reason. 4. Is the note hard to understand, incomplete, or not addressing the claim? Assign “notHelpfulHardToUnderstand” or “notHelpfulOffTopic” as required. 5. Would a typical, reasonably attentive reader gain essential, accurate insight into the claim’s truth, context, or credibility—including debunking of misused/incorrect supporting evidence—because of this note? If yes, “helpful” reasons most fitting the note’s substance. If no, most directly explanatory “notHelpful” reasons. Output in the following JSON format only, no extra output: {"helpfulness": helpful or non_helpful,"reasons":"reason1;reason2"}
CLAIM: \${claim}
NOTE: \${note}
Answer:
Table 16: Optimized English prompt for note helpfulness and reason prediction.
Seed Definitions
Helpful Reasons "helpfulClear": "**HelpfulClear** refers to the practice of validating and clarifying claims by providing accurate context or corrections through credible evidence and expert opinion. It involves critically examining information to dispel misinformation, address misconceptions, and ensure that communications are both accurate and informative. The aim is to enhance understanding, promote informed decision-making, and prevent the spread of false or misleading narratives.", "helpfulGoodSources": "**HelpfulGoodSources** refers to reliable, fact-based information that clarifies, corrects, or disproves misleading or false claims. It encompasses various forms of authoritative references, including government reports, scientific studies, expert opinions, and credible news articles. The purpose of these sources is to provide accuracy and context to statements made in public discourse, aiding in the dissemination of factual knowledge and supporting informed decision-making.", "helpfulAddressesClaim": "**HelpfulAddressesClaim** is a concept that refers to the clarification or correction of claims made in various contexts by providing accurate information, context, or details that counter or explain the original assertion. This often involves highlighting misconceptions, providing factual evidence, or referencing authoritative sources to support the revised understanding of the claim, ultimately helping to prevent misinformation and promoting informed discourse.", "helpfulImportantContext": "The REASON, defined as ḧelpfulImportantContext,r̈efers to the provision of essential background information that clarifies or corrects claims presented in various statements. This context is crucial for understanding the accuracy, validity, or implications of the claims made, by offering factual corrections, historical context, or relevant comparisons to ensure a more informed and nuanced comprehension of the situation or assertion being discussed.", "helpfulInformative": "The definition of the REASON ḧelpfulInformativeïs that it refers to explanations or notes provided in response to claims, which serve to clarify, correct inaccuracies, or provide essential context and background information. These notes help the reader better understand the claims by presenting factual evidence, linking to credible sources, and addressing misunderstandings, thereby enhancing the overall comprehension of the topic at hand.", "helpfulUnbiasedLanguage": "The definition of the REASON ḧelpfulUnbiasedLanguageïs the use of clear, neutral, and factual language that provides context and clarification to claims without introducing bias or misinformation. It seeks to present information in an objective manner, ensuring that the audience receives accurate and relevant details to better understand the claims being made. This approach emphasizes transparency and factual integrity, allowing for informed discussions and assessments.", "helpfulEmpathetic": "**Helpful Empathetic**: This reason reflects the intention to provide accurate and clarifying information in response to claims that are misleading or false. It embodies a commitment to enhancing understanding by offering factual corrections and context that reveal the truth, thereby enabling informed discussion. The goal is to support others in recognizing discrepancies or inaccuracies in claims, fostering a more informed perspective while acknowledging the importance of factual accuracy."
Unhelpful Reasons "notHelpfulOpinionSpeculation": "The definition of the REASON n̈otHelpfulOpinionSpeculationïs as follows: It refers to statements or notes that provide personal opinions, unverified claims, or speculative assertions that do not contribute factual information or clarification relevant to the claim being discussed. These opinions or speculative remarks often lack substantiation and do not advance the understanding of the subject matter, thus being non-helpful in the context of verifying or explaining the claims made.", "notHelpfulArgumentativeOrBiased": "The REASON "notHelpfulArgumentativeOrBiased" can be defined as follows: This classification refers to comments or notes that fail to provide constructive or relevant context to the accompanying claims. Instead, they tend to be argumentative, inflammatory, or biased, often lacking factual support or logical reasoning. Such responses may attempt to undermine the original claim without addressing its merits, focusing instead on personal attacks, irrelevant comparisons, or emotional appeals that detract from a rational discussion. In essence, they do not foster an understanding of the claim and instead perpetuate division or misinformation.", "notHelpfulMissingKeyPoints": "The definition of the REASON n̈otHelpfulMissingKeyPointsïs: The notes provided do not adequately address or clarify the claims made. Instead, they often introduce unrelated information or misconstrue the context, leaving significant gaps in understanding the primary arguments presented in the claims.", "notHelpfulNoteNotNeeded": "The definition of the REASON n̈otHelpfulNoteNotNeededïs that the provided notes fail to contribute meaningful or relevant information to clarify, support, or contradict the claims made. Instead, they either introduce unrelated details, emphasize inaccuracies without addressing the core of the claim, or illustrate a lack of factual backing, thus rendering them ineffective in providing substantiated context or understanding regarding the claims.", "notHelpfulIncorrect": "The definition of the REASON n̈otHelpfulIncorrectïs as follows: This category refers to claims that are supported by notes or explanations that are factually inaccurate, misleading, or irrelevant to the initial claim. The notes do not effectively clarify, substantiate, or provide accurate context for the claims, thereby failing to contribute meaningful information to the discussion. Instead, they may promote misinformation, misunderstandings, or misinterpretations of the original assertions.", "notHelpfulSourcesMissingOrUnreliable": "The REASON for the samples being non-helpful is defined as **the presence of sources that are unreliable, lack credibility, or do not substantiate the claims made**. This includes situations where evidence is anecdotal, opinion-based, or referencing non-verifiable documents, thereby failing to provide the necessary backing to support the claims.", "notHelpfulSpamHarassmentOrAbuse": "The definition of the REASON n̈otHelpfulSpamHarassmentOrAbuseïs content that is either unconstructive, irrelevant, or disrespectful in nature, often resorting to personal attacks, derogatory language, or baseless accusations. Such comments fail to provide meaningful engagement or contribute to the discussion, instead serving to provoke or demean individuals, typically lacking in factual or supportive content.", "notHelpfulIrrelevantSources": "The REASON n̈otHelpfulIrrelevantSourcesc̈an be defined as follows: NotHelpfulIrrelevantSources refers to a category of information where the claims being made are accompanied by explanations, notes, or sources that do not provide pertinent or factual support to validate the claims. This includes instances where the sources cited are either misleading, unrelated to the claim, or factually incorrect, resulting in a lack of credible evidence that would otherwise substantiate the assertions made. Essentially, information categorized under NotHelpfulIrrelevantSources fails to enhance understanding or clarify the claims due to its inaccuracy or irrelevance.", "notHelpfulOther": "The definition of the REASON n̈otHelpfulOtherïs that the provided notes do not effectively contribute to or clarify the claims made. Instead, they often reference unrelated information or distract from the central claim, failing to provide direct support, context, or factual information pertinent to understanding or verifying the claim.", "notHelpfulHardToUnderstand": "The REASON for the above samples being non-helpful is that the notes provided do not effectively clarify or support the claims made. They often lack relevant information, context, or clarity, making it difficult for readers to comprehend the claims or understand their significance. This results in a communication gap where the intended message is obscured or misunderstood.", "notHelpfulOpinionSpeculationOrBias": "The definition of the REASON n̈otHelpfulOpinionSpeculationOrBiasïs as follows: This reason applies to statements or claims that are influenced by subjective views, personal opinions, speculation, or perceived biases rather than presenting objective facts or useful information. Such claims often lack substantiation and do not contribute constructively to the discourse, focusing instead on emotional responses or distorted interpretations of events or behaviors."
Table 17: Note reason seed definitions.
Optimized Definitions
Helpful Reasons helpfulClear: The note delivers an accurate, unambiguous correction or direct clarification targeting a central claim or supporting fact/evidence. Use when the note dispels a major misconception, corrects a key number, identifies a misattributed element, or plainly states the essential fact needed to resolve or accurately interpret the claim. helpfulAddressesClaim: The note explicitly and accurately responds to or explains the principal assertion, central factual sub-claim, or foundational evidence of the claim. Includes refuting an implied mechanism or performing a core fact check that settles the claim’s credibility. helpfulGoodSources: The note cites sources that are authoritative, reputable, and faithfully, directly, and unambiguously support its factual correction or clarification. Pick only if sourcing is integral—don’t select if other reasons are more fundamental to helpfulness. helpfulImportantContext: The note supplies indispensable, targeted background or context needed for the reader to assess the claim or its evidence correctly (such as revealing regulatory nuance, historical policies, deception/fraud, or other context that transforms meaning or evaluation). helpfulInformative: The note gives essential factual explanation or contextualization directly linked to dispelling misunderstanding about the claim’s main point. Don’t use if the note merely provides general or background info. helpfulUnbiasedLanguage: The note’s language is clear, objective, and neutral while providing a substantive correction or clarification. Use only paired with a direct factual "helpful" label. helpfulEmpathetic: The note enhances the factual correction by presenting it in a manner that increases reader trust and comprehension through clarity and respect. helpfulUniqueContext: The note delivers information not widely available, central to correctly interpreting the claim or core evidence. Unhelpful Reasons notHelpfulOpinionSpeculation: The note is mainly opinion, speculation, or contains unsubstantiated assertions, not clarifying any core aspect of the claim or its evidence. notHelpfulArgumentativeOrBiased: Hostile, inflammatory, or overtly biased language dominates and no factual clarification or relevant context is given. notHelpfulMissingKeyPoints: The note wholly fails to clarify, correct, or address the main assertion, supporting fact, or evidence, OR it focuses solely on peripheral, partial, or unrelated points. notHelpfulNoteNotNeeded: The note adds no meaningful, fresh, or essential clarification—it is trivial, redundant, tangential, or consists of widely-known facts non-essential for comprehending the claim. notHelpfulIncorrect: Contains factual errors, misleading/unsupported claims, or confuses main assertions or evidence. notHelpfulSourcesMissingOrUnreliable: Sources are missing, non-credible, don’t support the note’s claim, or are misrepresented. notHelpfulSpamHarassmentOrAbuse: Contains spam, personal attacks, or offensive/unrelated content. notHelpfulIrrelevantSources: Sources in the note do not substantiate or clarify any main aspect of the claim or are not relevant to its core. notHelpfulOther: Any failure to clarify or support the claim not captured by the above categories. notHelpfulHardToUnderstand: The note is so unclear, incomplete, or poorly written that its meaning regarding the claim cannot be determined. notHelpfulOpinionSpeculationOrBias: Personal views, speculation, or bias present to such a degree that no clarification on the main claim or evidence is possible. notHelpfulOffTopic: The note is wholly unrelated to any claim feature or its main supporting facts/evidence.
Table 18: Optimized reason definitions.
Prompt for Fact Checking Climate Fever claims
Direct Prompt
Fact-check the following claim using provided evidence:
Claim: \${claim}
Evidence: \${evidence_text}
Classify the claim as SUPPORTS, REFUTES, NOT_ENOUGH_INFO or DISPUTED.
Format: Classification: [YOUR_ANSWER]
Brief reason:
Incorporating Helpfulness Information
Fact-check this claim using the evidence and the helpfulness information of the evidence, if the evidence is not helpful, take less weight of the evidence.
Claim: \${claim}
Evidence: \${evidence_text_with_helpfulness_information}
Classify the claim as SUPPORTS, REFUTES, NOT_ENOUGH_INFO or DISPUTED.
Format: Classification: [YOUR_ANSWER]
Brief reason:
Table 19: Prompt for Fact Checking CLIMATE-FEVER claims.